# SOUL: Unlocking the Power of Second-Order Optimization for LLM Unlearning

### v3

@@ Appendix A Additional Experimental Details and Results
Table A2 demonstrates the system prompts we incorporated before each question in the ToFU and copyright removal tasks, employing the input-based method as described by Thaker et al. (2024):
A.4 Unlearning configurations.
@@
A.5 Text-completion instructions.
Here are the text completion task instructions we used for the copyright removal evaluation. These prefixes are added before the original partial sentences from the Harry Potter books to guide the LLM in completing the texts.
@@
    Expand on this snippet, please:
A.6 Zero-shot tasks
For utility evaluation, we utilized the LM Evaluation Harness Gao et al. (2023) framework to assess the zero-shot accuracy across a suite of tasks, including BoolQ Clark et al. (2019), RTE Dagan et al. (2005), HellaSwag Zellers et al. (2019), Winogrande Sakaguchi et al. (2021), ARC-Challenge Chollet (2019), ARC-Easy Chollet (2019), OpenBookQA Mihaylov et al. (2018), and Piqa Bisk et al. (2020). The mean accuracy across these diverse tasks was computed and reported as a holistic measure of model utility post-unlearning.
A.7 Additional visualization
@@ F. Scott Fitzgerald is the author of ‘The Great Gatsby (1925)’
Examples for copyright removal
Table A5 provides examples of texts generated by unlearned LLaMA2-7B models subjected to various unlearning methods within the context of copyright removal tasks. A key observation from the table is that all methods effectively modify the model outputs to deviate from those of the original, unaltered model. However, instances persist where methods using first-order optimizers, such as FO-PO, produce content that bears relevance to Harry Potter, as exemplified by the mention of ‘Harry’ in the generated text from prompt 3. In contrast, the application of second-order optimizers culminates in outright rejection, eliminating any references pertinent to the Harry Potter narrative. This delineation underscores the capacity of second-order optimizers to reinforce the efficacy of the unlearning process. A similar phenomenon is also noted with the GradDiff method, further affirming the advantage of second-order optimization in achieving more thorough unlearning outcomes.
@@
Examples for LLMs detoxification task.

### v4

@@ B.1 Datasets, tasks and models
hyperparameters
@@
B.3 Evaluation metrics
To evaluate the effectiveness of fictitious unlearning in the TOFU task, we measure the distinguishability of statistical measures between the forget and retain sets using LLM-generated truthful ratios, as defined in the original TOFU benchmark Maini et al. (2024). This assessment is conducted via the Kolmogorov-Smirnov (KS) test. We utilize $1-$ p-value obtained from the KS test as the Forget Quality to assess unlearning effectiveness. In the experimentation, a high forget quality represents successful unlearning, indicating an increased distributional divergence between the forget and retain sets. We also measure unlearning effectiveness using the Membership Inference Attack (MIA) achieved through the Min-kProbability method Shi et al. (2023). This method determines whether a specific piece of text was part of an LLM’s training dataset. For our evaluation, we aim to detect the membership of the forgotten data as if it were part of the training set. We use data samples from world facts and real authors as the non-training test set and specifically measure the Area Under the Curve (AUC) of the Min-k%-based MIA detector in identifying whether the forgotten data was originally included in the training set. Ideally, a well-unlearned model should achieve a lower AUC, indicating improved unlearning effectiveness by not detecting forgotten data as part of the training set. Furthermore, we assess the unlearning performance of the LLM after unlearning (referred to as the unlearned model) by computing the Rouge-L recall against the ground truth and measuring the accuracy of the generated text. This involves comparing the cosine similarity of semantic embeddings from Sentence-BERT Reimers and Gurevych (2019) with both the ground truth and alternative incorrect responses in the TOFU dataset. Correctness is determined when the semantic embedding of the generated text is closest to the ground truth. We apply the same accuracy and Rouge-L recall metrics to evaluate utility preservation on sets related to retained information, real authors, and world facts.
In the copyright removal task, we randomly truncate 300 excerpts from the original Harry Potter dataset to the first $k$ tokens and evaluate them using BLEU and Rouge-L recall for prompt lengths of 100 and 300 tokens, with text completion instructions shown as following:
@@ B.3 Evaluation metrics
    Expand on this snippet, please:"
In the model detoxification task, toxicity is assessed using real toxic prompts Gehman et al. (2020) and the PKU-SafeRLHF test set Ji et al. (2024), assigning toxicity scores with Toxic-BERT Hanu and Unitary team (2020). For both the copyright removal and detoxification tasks, utility preservation is assessed using the LM Evaluation Harness Gao et al. (2023) to compute perplexity (PPL) on the Wikitext Merity et al. (2016) . We also assess the zero-shot accuracy across a suite of tasks, including BoolQ Clark et al. (2019), RTE Dagan et al. (2005), HellaSwag Zellers et al. (2019), Winogrande Sakaguchi et al. (2021), ARC-Challenge Chollet (2019), ARC-Easy Chollet (2019), OpenBookQA Mihaylov et al. (2018), and Piqa Bisk et al. (2020). The mean accuracy across these diverse tasks was computed and reported as a holistic measure of model utility post-unlearning. Additional evaluations include TruthfulQA Lin et al. (2021). Note that, similar to existing literature Eldan and Russinovich (2023); Maini et al. (2024), we did not consider more complex utility evaluations such as instruction-following ability. This is because the primary models are pre-trained LLMs not adapted using RLHF Achiam et al. (2023).
B.4 Additional visualization
@@ F. Scott Fitzgerald is the author of ‘The Great Gatsby (1925)’
‘The Great Gatsby’ was written by F. Scott Fitzgerald.
FO-NPO
‘The Great Gatsby’ was written by F. Scott Fitzgerald.
SO-NPO
@@
FO-NPO
The writer originating from Baku, Azerbaijan,
with a birth date of February 14, 1960, is Anara Yusifova..
SO-NPO
The writer originating from Baku, Azerbaijan,
with a birth date of February 14, 1960, is Anara Yusifova..
@@
Examples for copyright removal
Table A5 provides examples of texts generated by unlearned LLaMA2-7B-chat models subjected to various unlearning methods within the context of copyright removal tasks. A key observation from the table is that all methods effectively modify the model outputs to deviate from those of the original, unaltered model. However, instances persist where methods using first-order optimizers, such as FO-PO, produce content that bears relevance to Harry Potter, as exemplified by the mention of ‘Harry’ in the generated text from prompt 3. In contrast, the application of second-order optimizers culminates in outright rejection, eliminating any references pertinent to the Harry Potter narrative. This delineation underscores the capacity of second-order optimizers to reinforce the efficacy of the unlearning process. A similar phenomenon is also noted with the GradDiff method, further affirming the advantage of second-order optimization in achieving more thorough unlearning outcomes.
@@
Method Forget efficacy Utility
Real Toxicity Prompts PKU-SafeRLHF PPL Zero-shot Acc. TruthfulQA
Toxic Score Toxic Score
OPT-350M
Original 0.0833 0.1166 25.43 42.69% 0.2387
FO-GradDiff 0.0744 0.1048 26.30 43.36% 0.2313
SO-GradDiff (Ours) 0.0737 0.0555 26.78 43.29% 0.2289
FO-PO 0.0491 0.0460 26.11 42.39% 0.2411
SO-PO (Ours) 0.0424 0.0356 26.20 43.08% 0.2448
OPT-1.3B
Original 0.0807 0.1118 16.49 48.16% 0.2411
FO-GradDiff 0.0748 0.0673 30.87 41.16% 0.2362
SO-GradDiff (Ours) 0.0561 0.0618 28.77 40.34% 0.2240
FO-PO 0.0404 0.0253 18.26 46.25% 0.2852
SO-PO (Ours) 0.0335 0.0165 17.97 48.60% 0.2742
LLaMA2-7B
Original 0.0710 0.1027 8.79 62.08% 0.2521
FO-GradDiff 0.0708 0.0989 8.77 61.38% 0.2534
SO-GradDiff (Ours) 0.0722 0.0987 8.79 61.32% 0.2534
FO-PO 0.0626 0.0790 8.78 61.92% 0.2632
SO-PO (Ours) 0.0528 0.0443 8.87 62.80% 0.2656
Table A7: Performance comparison between SOUL and its FO counterparts in the task of model detoxification, following the format of Table 4.
Examples for LLMs detoxification task.
B.5 Results on LLM detoxification
In Table A7, we demonstrate that the proposed SO unlearning methods effectively reduce the toxicity score on both the Real Toxicity Prompts and PKU-SafeRLHF datasets while maintaining or even improving utility. For instance, in the LLaMA2-7B model, SO-PO achieved a clear reduction in the toxic score on the PKU-SafeRLHF dataset and showed enhanced performance in zero-shot accuracy compared to FO-PO. This indicates improved unlearning efficacy of SOUL without sacrificing model utility. In addition, Table A6 includes visualizations that exemplify the outputs after the application of unlearning to the LLaMA2-7B models. These visualizations further corroborate that SO optimizers improve unlearning efficacy, particularly highlighting that SO-PO achieves the most effective unlearning performance.
B.6 Performance comparison between IU and SOUL
In this section, we compare the performance of SOUL with that of traditional influence unlearning Izzo et al. (2021); Koh and Liang (2017) in Table A8. This comparison demonstrates that merely adapting IU for LLM unlearning does not yield satisfactory unlearning effectiveness due to its static nature and lack of optimization power. However, SOUL improves upon this by transitioning from the static, one-shot nature of influence unlearning to an iterative, optimization-driven influence-aware approach.
Method Unlearning Efficacy Utility
Forget Retain Real Authors World Facts
Forget quality Acc. Rouge-L MIA Acc. Rouge-L Acc. Rouge-L Acc. Rouge-L
Original 0.36 85.25% 0.9796 0.7894 85.75% 0.9825 89.00% 0.9330 86.32% 0.8960
IU 0.36 84.25% 0.9573 0.7881 86.00% 0.9414 85.00% 0.9390 83.76% 0.8746
SOUL 1.00 10.25% 0.0221 0.2156 72.25% 0.5960 78.00% 0.8113 82.05% 0.8675
Table A8: Performance comparison between SOUL and IU (influence unlearning), following the format of Table 2.
B.7 Adversarial evaluation for SOUL
Methods Forget acc. $\downarrow$
Forget acc. $\downarrow$
FO-GradDiff 72.25% 72.25%
SO-GradDiff 10.25% 16.00%
FO-PO 37.00% 37.00%
SO-PO 28.75% 31.25%
FO-NPO 16.00% 25.00%
SO-NPO 16.00% 20.00%
Table A9: Forget accuracy in the absence or presence of jailbreak prompt for different unlearning methods on the TOFU dataset.
Furthermore, we evaluate the unlearning effectiveness in the presence of jailbreak prompts, generated following the method in Lynch et al. (2024). This assesses whether the forgotten knowledge can be revoked when tested using a jailbreak prompt, such as a question-answer pair from the retain set that enforces non-forgetting. Note that this can be regarded as a non-optimization based jailbreaking attack for LLMs post-unlearning. Table A9 presents the forget accuracy comparisons before and after jailbreaking across different unlearning methods. While jailbreaking could degrade unlearning efficacy (as evidenced by the increase in forget accuracy), SOUL consistently achieves lower forget accuracy compared to first-order methods after jailbreaking. This indicates the robustness benefit of using SOUL. In addition, since the design of jailbreak prompts in Lynch et al. (2024) is not based on an optimization approach, these prompts may become ineffective at attacking LLMs post-unlearning, as evidenced by the same forget accuracy after jailbreaking.
B.8 Time analysis
Methods
Running Time
FO-GradDiff 30
SO-GradDiff 30
FO-PO 30
SO-PO 31
FO-NPO 32
SO-NPO 35
Table A10: Time comparison among different methods on the TOFU task.
In our experiments, we configured the Hessian updating frequency in Sophia Liu et al. (2023a) to update the Hessian at each optimization step. Despite the potential for high computational demand, this approach remains computationally efficient because Sophia approximates the diagonal values of the Hessian using the element-wise square of the gradient. This approximation significantly reduces the additional computational overhead, making it minimal compared to exclusive reliance on first-order updates. Table A10 presents the running time costs for various methods applied to the TOFU task, demonstrating that the use of a second-order optimizer does not incur a significantly greater overhead compared to methods that employ first-order optimizers.